OpenStudy (anonymous):

Find an equation of the line through (3,7) and parallel to y=3x+2. Write in function notation.

OpenStudy (accessdenied):

Did you have any ideas on this problem so far? Is there a particular part that is causing trouble?

OpenStudy (anonymous):

I'd like to know how to set it up for solving.

OpenStudy (accessdenied):

We know the equation is of a line. We also know that the line we want is parallel to the line y = 3x + 2. So that gives us information about the slope, right? Combined with the point we are given, we can use point-slope form to set up the equation.

OpenStudy (anonymous):

What's the point-slope formula again? Pardon, I'm a bit fried tonight.

OpenStudy (accessdenied):

\( y - y_0 = m(x - x_0) \) Where \( (x_0, y_0 ) \) is your given point. Although once we find this equation, I think we want to change out y = f(x) for the function notation.

OpenStudy (anonymous):

So, y-7=3(x-3)?

OpenStudy (anonymous):

Then distribute?

OpenStudy (accessdenied):

Yup, that looks good.

OpenStudy (accessdenied):

And yes you could distribute. Then add the 7 to both sides and you'd have your line equation. y = f(x) would finish off the problem

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (accessdenied):

You're welcome! :)